# Prodiabatic Elimination: Higher Order Elimination of Fast Variables with Quantum Noise

### v1

@@ Abstract
Introduction– Exact descriptions of open quantum dynamics are rarely available, making approximation methods essential for constructing effective models. Systems exhibiting well-separated time scales are particularly amenable to such approaches, as the separation can be exploited to derive simplified effective models [6], with applications in physics [34, 11, 2], chemistry [20], biology [8], medicine [22], and even finance [19]. One of the most prominent approaches to tackle systems with a separation of timescales is provided by the adiabatic elimination [17, 34, 28, 26, 11, 15, 38, 35], in which fast degrees of freedom are systematically removed, yielding an effective description for the remaining slow variables.
@@
The prodiabatic elimination– We model the dynamics of the cavity field with the quantum Langevin equation [16] ($\hbar=1$)
@@
The conditions, apart from $\Omega/\kappa$, are known as the bad-cavity limit [34], implying the Purcell factor $F_{\rm p}=\frac{4g^{2}}{\gamma\kappa}$ to be of order $1$. The last condition in Eq. (3) may be relaxed to $\mathcal{O}\left(f/\kappa\right)=1$, see [1].
@@
The leading-order adiabatic elimination can be obtained by setting $\dot{\hat{a}}=0$ in Eq. (1) and solving for $\hat{a}$. Dropping the noise term, which is justified for time- and normal-ordered expectation values [34], results in $\hat{a}=\hat{a}_{\rm adb}=-\frac{2it_{\rm c}}{\kappa}[g\hat{b}(t)-f(t)]$,
where $t_{\rm c}=\left(1+\frac{2i\Delta}{\kappa}\right)^{-1}$ is the cavity susceptibility. Note that $\hat{a}_{\rm adb}$ is an atomic operator, it does not act on the cavity. The correlation functions in Eq. (4) may then be evaluated using the quantum regression theorem [24, 10], see also Ref. [1].
@@
Jaynes-Cummings model–As a first example, we consider a two-level system as the atom with $|\hat{b}=\hat{\sigma}\equiv\outerproduct{0}{1}$ and $\hat{v}=\hat{\sigma}$. Eq. (2) then reduces to
@@
Figure 3: Prodiabatic elimination for STIRAP protocol. Semi-transparent: numerical simulations, dashed: adiabatic elimination, solid: prodiabatic elimination and doted: Lindblad master equation Eq. (S34). The drives are boxcar functions: $f_{\rm V}(t)=\kappa$ if $|55-\kappa t|\leq 10$ and $f_{\rm H}(t)=\kappa$ if $|45-\kappa t|\leq 10$, further $g/\kappa=1/10$ and $\gamma/\kappa=5\cdot 10^{-4}$.
@@ 11 Place Marcelin Berthelot, 75321 Paris Cedex 05, France
In this supplementary material, we give additional formulas and derivations supporting the statements of the main text including: the general form of the Lindblad master equation our approach is applicable to, the derivation of Eq. (11), the use of the quantum regression theorem for multi-time correlation functions, the Lindblad master equations equivalent to the dynamics of the atom after the prodiabatic elimination for both examples considered, a further exploration of parameter space for STIRAP, and extending the approach to stronger drives.
@@
Equations (12) can be obtained from the general quantum Langevin equation in the presence of two input channels,

### v2

# arXiv:2602.21896v2[quant-ph] 02 Mar 2026

@@ Abstract
Introduction.—Exact descriptions of open quantum dynamics are rarely available, making approximation methods essential for constructing effective models. Systems exhibiting well-separated time scales are particularly amenable to such approaches, as the separation can be exploited to derive simplified effective models [6], with applications in physics [34, 11, 2], chemistry [20], biology [8], medicine [22], and even finance [19]. One of the most prominent approaches to tackle systems with a separation of timescales is provided by the adiabatic elimination [17, 34, 28, 26, 11, 15, 38, 35], in which fast degrees of freedom are systematically removed, yielding an effective description for the remaining slow variables.
@@
The prodiabatic elimination.—We model the dynamics of the cavity field with the quantum Langevin equation [16] ($\hbar=1$)
@@
The conditions, apart from $\Omega/\kappa$, are known as the bad-cavity limit [34], implying the Purcell factor $F_{\rm p}=\frac{4g^{2}}{\gamma\kappa}$ to be of order $1$.
@@
The leading-order adiabatic elimination can be obtained by setting $\dot{\hat{a}}=0$ in Eq. (1) and solving for $\hat{a}$. Dropping the noise term, which is justified for time- and normal-ordered expectation values [34], results in $\hat{a}=\hat{a}_{\rm adb}=-\frac{2it_{\rm c}}{\kappa}[g\hat{b}(t)-f(t)]$, where $t_{\rm c}=\left(1+\frac{2i\Delta}{\kappa}\right)^{-1}$ is the cavity susceptibility. Note that $\hat{a}_{\rm adb}$ is an atomic operator, it does not act on the cavity. The correlation functions in Eq. (4) may then be evaluated using the quantum regression theorem [24, 10].
@@
Jaynes-Cummings model.—As a first example, we consider a two-level system as the atom with $|\hat{b}=\hat{\sigma}\equiv\outerproduct{0}{1}$ and $\hat{v}=\hat{\sigma}$. Eq. (2) then reduces to
@@
Figure 3: Prodiabatic elimination for STIRAP protocol. Semi-transparent: numerical simulations, dashed: adiabatic elimination, solid: prodiabatic elimination and doted: Lindblad master equation [1]. The drives are boxcar functions: $f_{\rm V}(t)=\kappa$ if $|55-\kappa t|\leq 10$ and $f_{\rm H}(t)=\kappa$ if $|45-\kappa t|\leq 10$, further $g/\kappa=1/10$ and $\gamma/\kappa=5\cdot 10^{-4}$.
@@
Acknowledgements.
Acknowledgments.—This work was supported by the Swiss National Science Foundation (Eccellenza Professorial Fellowship PCEFP2_194268). J.N. acknowledges support from the Eurpean Research Council (ERC Project ’Cocoquest’ 101043705). M.B. acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 101002955 – CONQUER).
@@ 11 Place Marcelin Berthelot, 75321 Paris Cedex 05, France
In this supplementary material, we give additional formulas and derivations supporting the statements of the main text including: the general form of the Lindblad master equation our approach is applicable to, the derivation of Eq. (11), the Lindblad master equations equivalent to the dynamics of the atom after the prodiabatic elimination for both examples considered and a further exploration of parameter space for STIRAP.
@@
Equations (1, 2) can be obtained from the general quantum Langevin equation in the presence of two input channels,